The following table lists the components needed to assemble an end item, lead times, and quantities on hand.

Item End B C D E F G H
LT (wk) 2 2 2 2 1 2 1 2
Amount on hand 0 11 8 26 12 28 6 0

End Item B (2) E (2) F(3) G (2) E (2) H14) E (2)

a. If 20 units of the end item are to be assembled, how many additional units of E are needed? (Hint: You don't need to develop an MRP plan to determine this.)

Additional units              

b. An order for the end item is scheduled to be shipped at the start of week 17. What is the latest week that the order can be started and still be ready to ship on time? (Hint: You don't need to develop an MRP plan for this part either.)

The latest week             
Answer #1

Answer a: 138 additional units of E are needed

Explanation:

Item Relation Requirement= relation *net requirement of related part Inventory net requirement= requirement- Inventory
End, A 20 0 20
B 2A 40 11 29
C 1A 20 8 12
D 3A 60 26 34
E 2B +2C+2D 150 12 138

Answer b: the latest week to start is 11

Explanation:

Start time= 17
BOM path Total lead time= sum of all lead times in a path Start time= Start week -total lead time of path
ABE 2+2+1= 5 17-5= 12
ABF 2+2+2= 6 17-6= 11
ACG 5 12
ACE 5 12
ADH 6 11
ADE 5 12

we have to start by 11th week.

Please ask, if you have any doubts through the comment section. Do rate the answer